OpenStudy (anonymous):

Prove \[ \sum_{n=1}^\infty \frac { n(n-1)}{2^n} =4 \] Hint: Use \[ \sum_{n=0}^\infty x^n =\frac 1 {1-x}, \quad |x|<1 \] and use the term by term differentiation method. Of course, I know how to do it. I am sure some of you can. Give it a try.

OpenStudy (anonymous):

The problem needs differentiation twice. This one needs three: \[ \sum_{n=3}^\infty \frac { n(n-1)(n-2)}{2^n} =12 \]

OpenStudy (anonymous):

Why differentiation? It isn't continuous is it?

OpenStudy (anonymous):

for |x| < 1 it is continuous and differentiable.

OpenStudy (anonymous):

Hint: Use \[ \sum_{n=0}^\infty x^n =\frac 1 {1-x}, \quad |x|<1 \] Differentiate both sides to get \[ \sum_{n=1}^\infty n x^{n-1} =\frac {-1} {(1-x)^2}, \quad |x|<1 \] Differentiate another time and multiply both sides by \[x^2\] and you can find the first sum.

OpenStudy (experimentx):

one question: why change from n=0 to n=1 on differentiation??

OpenStudy (anonymous):

Yeah, I should have considered your hint earlier :(

OpenStudy (anonymous):

1st is Zero anyway

OpenStudy (anonymous):

I mean for n=0

OpenStudy (experimentx):

Oo.. i get it.

OpenStudy (anonymous):

Yes, you did.

OpenStudy (experimentx):

for x=(1/2) we get, 2*(x)^2/(1-x)^3 = 4 , on the left hand side summation n=2 to inf n(n-1)(1/2)^2 = 4

OpenStudy (experimentx):

Problem was half solved by you anyway :D

OpenStudy (anonymous):

@experimentX do the problem that involves 3 differentiations

OpenStudy (experimentx):

i'll give a try

OpenStudy (anonymous):

You are almost there.

OpenStudy (experimentx):

yeah i get it, multiply both sides by x^3 and replace x=1/2, we get the required expression on left hand side, on the right hand side, we have \[\frac{6x^3}{(1-x)^4} = \frac{6*(1/2)^3}{(1/2)^4} = 12 \]

OpenStudy (anonymous):

Excellent. What is your background in Math?

OpenStudy (experimentx):

Physics major

OpenStudy (anonymous):

From my experience, Physics majors are very good in Math,

OpenStudy (experimentx):

I am quite not ... still struggling with calculus, though i appeared math major with algebra, analysis, advanced calculus, and mechanics ... i never understood anything .. esp algebra and analysis.

OpenStudy (anonymous):

You will get there. Practice, Practice and Practice.

OpenStudy (experimentx):

therefore still struggling on physics ... at least i hope to finish differential equations and multivariable calculus ...

OpenStudy (anonymous):

Use also my site http://saab.org to practice math from algebra to calculus to actuarial problems.

OpenStudy (anonymous):

I will post soon another seires problem. Try to do it. I will close this problem for now.

OpenStudy (experimentx):

thanks .. i will. you don't look like a student ... are you teacher at some university?